OpenStudy (anonymous):

f(x,y)=(x^2+y^2) e^-x find the rate of change of f at (1,2) as you move in the direction <-2,3>

OpenStudy (zarkon):

where are you stuck?

OpenStudy (anonymous):

i found the derivatives and I know i have to multiply by <-2/sqrt15, 3/sqrt15> i am just not sure what i do with the (1,2) if i put it in to the function or?

OpenStudy (zarkon):

make a unit vector

OpenStudy (zarkon):

\[\left<f_x(a,b),f_y(a,b)\right>\cdot\frac{<a,b>}{\|<a,b>\|}\]

OpenStudy (anonymous):

so is it (-2e^-x)/sqrt(13)

OpenStudy (zarkon):

as a final answer?

OpenStudy (zarkon):

there should be no x's in your answer

OpenStudy (zarkon):

your answer should just be a number

OpenStudy (anonymous):

so (-2sqrt(13))/(13e)

OpenStudy (zarkon):

no

OpenStudy (zarkon):

where are you getting the \(\sqrt{13}\)?

OpenStudy (zarkon):

oops...just a sec...I was using the wrong number

OpenStudy (zarkon):

i got \[\frac{6\sqrt{13}}{13e}\]

OpenStudy (zarkon):

my formula is a little off too.. \[\left<f_x(a,b),f_y(a,b)\right>\cdot\frac{<\alpha,\beta>}{\|<\alpha,\beta>\|}\] would be better

OpenStudy (anonymous):

what did you get for the partial derivative of x

OpenStudy (zarkon):

\[(-x^2+2x-y^2)e^{-x}\]

OpenStudy (zarkon):

\[\frac{\partial}{\partial x}(x^2+y^2) e^{-x}=(x^2+y^2)e^{-x}(-1)+(2x)e^{-x}\] \[=(-x^2-y^2)e^{-x}+(2x)e^{-x}=(-x^2+2x-y^2)e^{-x}\]

OpenStudy (anonymous):

okay thanks i see what i did wrong

OpenStudy (zarkon):

good :)

OpenStudy (anonymous):

thank you so much

OpenStudy (zarkon):

no problem